OpenStudy (anonymous):

Please help 2 Questions Bellow

OpenStudy (anonymous):

Here is question

OpenStudy (superdavesuper):

Very strange way to express the inequality. first is to write out the inequality shown on the number line first. can u do that?

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Sorry i dont know how to do that

OpenStudy (superdavesuper):

fill in the blank here plz: y ____ -7 ok?

OpenStudy (anonymous):

I will try

OpenStudy (anonymous):

Fill in the blank with a ≥or ≤ key

OpenStudy (superdavesuper):

\[> \ge \le \]

OpenStudy (superdavesuper):

\[> \ge \le < \]

OpenStudy (anonymous):

can you draw and show how you do it

OpenStudy (superdavesuper):

No ur question already has the line, u just need to read it and understand what it means.

OpenStudy (anonymous):

Yeah its -7

OpenStudy (superdavesuper):

OK but what is the sign that goes between y and -7?

OpenStudy (anonymous):

B?

OpenStudy (anonymous):

I mean a

OpenStudy (superdavesuper):

That is coreect :) Noooooooo why u change ur mind??

OpenStudy (anonymous):

Oh lol

OpenStudy (anonymous):

But how is that the answer

OpenStudy (superdavesuper):

Well actually im not sure - wat does the line w/ a circle mean? is it > or \[ \ge \]

OpenStudy (anonymous):

I don't know

OpenStudy (superdavesuper):

Looks like u are right afterall - the circle mean > so A is the right answer :)

OpenStudy (anonymous):

1 more question

OpenStudy (anonymous):

Yay

OpenStudy (superdavesuper):

multiply left n right by 2 first

OpenStudy (anonymous):

what do you mean left n

OpenStudy (superdavesuper):

left n right hand side of the inequality

OpenStudy (anonymous):

thank you

OpenStudy (superdavesuper):

Got it already? good job :)